John Fare purchased $18,000 worth of equipment by making a $3000 down payment and promising to pay the remainder of the cost in semiannual payments over the next 5 years. The interest rate on the debt is 6%, compounded semiannually. Find the following. (Round your answers to the nearest cent.) (a) the size of each payment (b) the total amount paid over the life of the loan (c) the total interest paid over the life of the loan
Answer #1

PV =18000
Amount of Loan =18000-3000 =15000
Number of Periods =5*2=10
Rate per semiannual =6%/2 =3%
a) Size of each Payment =PV/((1-(1+r)^-n)/r) =15000/((1-(1+3%)^-10)/3%) =1758.4576

b) Total amount paid =Number of Periods* Size of each payment =10*1758.4576 = 17584.58

c) Total interest paid over the life =17584.58-15000=2584.58